OpenStudy (anonymous):

Check answer Please and help with #11. Suppose houses are available for purchase during construction of a new housing development. Two different types of houses are being built: rancher and colonial. The mean price of the ranchers is $319,000 with a standard deviation of $12,000. The mean price of the colonials is $405,000 with a standard deviation of $18,500. 8. Find the combined mean price of the two different types of houses. (2 points) $362,000 9. Find the combined standard deviation of the two different types of houses. (2 points) $22,051.0771 or $22,051.08 10. The purchase price for the colonial houses is reduced by $10,000. What would be the new mean for the colonial houses? (2 points) $357,000 11. If the purchase price for the colonial houses is reduced by $10,000, how much would the standard deviation for the colonial houses change? (2 points)16,055.72 or changed to $5,996.09

OpenStudy (anonymous):

A) Combined mean = (319000+405000)/2 = 362,000 B) Combined variance = (12000)^2+(18500)^2 =144 000 000 + 342 250 000 = 486 250 000 Combined standard deviation = square root [486 250 000] = 22,051.0771 C) colonial new price = 395,000 new mean = (319,000+395,000)/2 = D) Do as in (B) using the prices 319,000 and 395,000

OpenStudy (anonymous):

Is the answer for correct and which one.

OpenStudy (anonymous):

first

OpenStudy (anonymous):

11. 16,055.72